OpenStudy (anonymous):

I WILL FAN AND MEDAL. What is the quotient in simplified form? State any restrictions on the variable. --> g-6 over g+2 / g+4 over g^2-g-6

OpenStudy (anonymous):

@Abhisar @ganeshie8 @Nnesha Any help ? Didn't understand at all

OpenStudy (anonymous):

@pooja195 @nincompoop

Nnesha (nnesha):

\[\huge\rm \frac{ \frac{ g- 6}{ g+2 } }{ \frac{ g+4 }{\color{red}{ g^2-g-6}} }\] like this ?

OpenStudy (anonymous):

nope, put the g+4 over g^2-g-6 at the right with the division sign in the middle @Nnesha

Nnesha (nnesha):

same thing :=)

Nnesha (nnesha):

\[\huge\rm \frac{ \frac{ g- 6}{ g+2 } }{ \frac{ g+4 }{\color{red}{ g^2-g-6}} }=\frac{g-6}{g+2}\div\frac{g+4}{g^2-g+6}\]

Nnesha (nnesha):

1st) try to factor out the equation g^2-g-6 how would you find factors of quadratic equation ?

OpenStudy (anonymous):

i dont know tbh @Nnesha

Nnesha (nnesha):

find two number when you multiply them you should get product of AC when you add or subtract the same two numbers you should get middle term a ,c and middle term is \[\rm Ax^2+Bx+C\]quadratic equation where a= leading coefficient b=middle term c=constant so what is a b and c in this equation g^2-g+6 ?

OpenStudy (anonymous):

a=1. b=1 and c=-6 ?

Nnesha (nnesha):

b=-1 b is negative so ac = -6 what two numbers you should multiply to get -6 but when you `add` them you should get -1

OpenStudy (anonymous):

-3 and 2

Nnesha (nnesha):

yes right so (g-3)(g+2) are factors of that equation \[\huge\rm \frac{g-6}{g+2}\div\frac{g+4}{(g-3)(g+2)}\] now change division to multiplication multiply first fraction with the `reciprocal` of the 2nd fraction

Nnesha (nnesha):

i'll give u an example

Nnesha (nnesha):

\[\huge\rm \frac{ a }{ b } \div \frac{ d }{ c }=\frac{ a }{ b } \times \frac{ c }{ d }\]

OpenStudy (anonymous):

i give up i dont understand haha i just need the answer tbh

Nnesha (nnesha):

it's not that hard just flip the 2nd fraction and then change division sign to multiplication that's it

Nnesha (nnesha):

if i have 2/3 then its reciprocal would be 3/2